# Beyond directed hypergraphs: heterogeneous hypergraphs and spectral centralities

### v1

@@ 4.2.1 $\mathcal{B}$-hypergraphs
This makes it perfectly suitable for the HEC centrality calculations, as the free index in the sum is the first one, and the sums run equally over the remaining indices. It is important to notice that the $\mathcal{H}$-eigenvector one needs to compute is that corresponding to the transposed hypergraph, which will be the one the one whose connectivity is called into question.
@@
Definition 4.4 ($\mathcal{B}$-uniformity)
@@
We will discuss and compute numerically centralities of hypergraphs fulfilling this constraint, but before that let us examine the other kind of hyperedges.
@@ 4.2.2 $\mathcal{F}$-hypergraphs
We now turn to the opposite case, that of directed hypergraphs containing only $\mathcal{F}$-hyperedges, with a fixed number of nodes per hyperedge (i.e. $m$-uniform)333 This type is employed in [22, 14] with a specifically weighted adjacency tensor . As we will see, this case is rather troublesome when it comes to centrality computations.
@@ 4.2.3 General directed hypergraphs
Definition 4.5 (HEC of a directed hypergraph)
@@
Although this definition may seem daunting, operationally it is rather clear: Given a $\mathcal{B}$-uniform hypergraph where $m_{B}$ is the number of head nodes (fixed due to the uniformity constraint), we turn every hyperedge into several ones, one per tail node, adjusting the combinatorial factors to avoid overcount due to the symmetry of the adjacency tensor. We are then left with a tensor $\mathcal{T}\in\mathbb{R}^{[m_{B},N]}$, whose centrality we can compute, as in the undirected case [9], as its Perron-like eigenvector.
@@ 4.2.4 Numerical examples
Finding datasets of directed hypergraphs freely available is far from trivial, especially if we are interested in real data ones rather than synthetic ones. However, we managed to construct several of them related to chemical [23] and astro-chemical reactions [^4] We also constructed others from metabolical reactions [19], but for the sake of conciseness the computation of the rankings and comparisons are left in the repository, see Data Availability section. [18]. Once the hypergraphs have been constructed, we find the strongly connected sub-hypergraph satisfying the $\mathcal{B}$-uniformity condition, if any.
From these hypergraphs we can then compute their directed $\mathcal{H}$-eigenvector centrality (using the code we already developed and made available in [16]) and that of their projection graph. The top $5$ nodes in each example, as obtained via the standard eigenvector centrality of the projected graph, as well as with the newly defined measure for directed hypergraphs, are shown for comparison in Table 1.
@@ 5 HCO HCO He C+ NO HNO CN H+
We can also compare these measures in terms of the Spearman’s $\rho$ coefficient between both measures, which measures the similarity of two rankings (ranging from 1, meaning identical rankings, to -1, meaning completely opposite rankings). The correlation between rankings for the first $K$ components of each hypergraph are shown in Figure 3.
@@
From these figures we can see that the rankings from the pairwise projection networks and the ones from our new method are positively correlated (as they should), being very similar in some cases (e.g. subfigures a and b) but different enough that they represent a different meaning of what importance is.
@@ 4.3 $k$-step hypergraphs
In [21] the authors designed a new centrality measure (the two-step eigenvector centrality) for undirected graphs, which consisted of constructing a tensor (the two-step tensor) from powers of the adjacency matrix of the graph, which could then leverage the machinery of $\mathcal{H}$-eigenvectors in order to rank the nodes of the original graph [^5] Said article is only concerned with undirected graphs and $\mathcal{H}$-eigenvectors, however there is actually no roadblock for extending it to directed graphs and $\mathcal{Z}$-eigenvectors..
Definition 4.6 (2-step centrality of a graph [21])
@@
Definition 4.7 ($k$-step hyperedge)
@@
Clearly, for an undirected network we have the symmetry property $T_{i_{1}i_{2}\dots i_{k-1}i_{k}}=T_{i_{k}i_{k-1}\dots i_{2}i_{1}}$, which is why in [21] they safely ignore the transposition step, although omitting this discussion.
@@
In some sense this measures provides centrality to the target node from its $k-1$ ancestors. This is very reminiscent of centrality measures involving the iterated line graph or the Hashimoto matrix in standard networks [24, 25].
@@
Theorem 4.8 (Strong connectedness of the hypergraph induced by $T_{i_{1}\dots i_{k}}$)
@@
Proof: Consider the definition of strong connectivity based on the graph $G^{M}$ (see Definition 2.4)
@@
because if $G$ is strongly connected then there must be at least one $(k-2)$-long walk starting from $j$ (possibly repeating edges and nodes). Therefore $M\geq A$ component-wise, so if $G$ is strongly connected, $G^{M}$ is strongly connected and so is $H$. $\square$
@@
In the last subfigure, corresponding to a flight transportation network, we find it remarkable that the rankings are almost identical, having the Spearman correlation very close to 1 for all $K$ sampled. This last dataset corresponds to an undirected network, which, for the purposes of this articles, is treated as a directed one with both directions per edge. However this bears no relation with the high correlation, and indeed we have checked that other undirected networks (e.g. the US Power Grid) do not have this feature.
@@ 5 Conclusions
Our paper sheds light on the wide variety of non-undirected hyperedges, beyond traditional directedness, acknowledging diverse forms of multiple interactions among individuals. By introducing the term “heterogeneous hyperedge”, we aimed to broaden the understanding of non-undirected interactions, including directed hyperedges as a specific type.
@@
Acknowledgements
This work has been partially supported by projects M3033 (URJC Grant) and the NextGeneration-EU funds. G. C-A. is funded by the URJC fellowship PREDOC-21-026-2164.
@@ Acknowledgements
The data used for the numerical results presented in this work, as well as the code written to analyze and plot them, can be found in the repository https://github.com/LaComarca-Lab/DirectedHyperCentrality.

### v2

@@ 4.2.1 $\mathcal{B}$-hypergraphs
This makes it perfectly suitable for the HEC centrality calculations, as the free index in the sum is the first one, and the sums run equally over the remaining indices. It is important to notice that the $\mathcal{H}$-eigenvector one needs to compute is that corresponding to the transposed hypergraph, which will be the one whose connectivity is called into question.
@@
Definition 4.4 ($\mathcal{B}$-uniformity).
@@
We will discuss and compute numerically centralities of hypergraphs fulfilling this constraint in a later subsection, but before that let us examine the other kind of hyperedges.
@@ 4.2.2 $\mathcal{F}$-hypergraphs
We now turn to the opposite case, that of directed hypergraphs containing only $\mathcal{F}$-hyperedges, with a fixed number of nodes per hyperedge (i.e. $m$-uniform)444 This type is employed in [xie2016spectral, qi2017tensor] with a specifically weighted adjacency tensor . As we will see, this case is rather troublesome when it comes to centrality computations.
@@ 4.2.3 General directed hypergraphs
Definition 4.5 (HEC of a directed hypergraph).
@@
Although this definition may seem daunting, operationally it is rather clear: Given a $\mathcal{B}$-uniform hypergraph where $m_{B}$ is the number of head nodes (fixed due to the uniformity constraint), we turn every hyperedge into several ones, one per tail node, adjusting the combinatorial factors to avoid overcount due to the symmetry of the adjacency tensor. We are then left with a tensor $\mathcal{T}\in\mathbb{R}^{[m_{B},N]}$, whose centrality we can compute as its Perron-like eigenvector, like in the undirected case [benson2019three].
@@ 4.2.4 Numerical examples
Finding datasets of directed hypergraphs freely available is far from trivial, especially if we are interested in real data ones rather than synthetic ones. However, we managed to construct several of them related to chemical [RTG-datasets] and astro-chemical reactions [^5] We also constructed others from metabolical reactions [BiGG-datasets], but for the sake of conciseness the computation of the rankings and comparisons are left in the repository, see Data Availability section. [KIDA-datasets]. Once the hypergraphs have been constructed [XGI-library], we find the strongly connected sub-hypergraph satisfying the $\mathcal{B}$-uniformity condition, if any.
From these hypergraphs we can then compute their directed $\mathcal{H}$-eigenvector centrality (using the code we already developed and made available in [contrerasaso2023uplifting]) and that of their projection graph. The top $5$ nodes in each example, as obtained via the standard eigenvector centrality of the projected graph, as well as with the newly defined measure for directed hypergraphs, are shown for comparison in Table 1.
@@ 5 HCO HCO He C+ NO HNO CN H+
We can also compare these measures in terms of the Spearman’s $\rho$ coefficient between both measures, which measures the similarity of two rankings (ranging from 1, meaning identical rankings, to -1, meaning completely opposite rankings). The correlation between rankings for the first $K$ components of each hypergraph are shown in Figure 3. Notice that, even though the covariance is symmetric, the rankings H3-E and E-H3 differ: that is due to the fact that, before the last datapoint, the correlation compares the first $K$ nodes in each measure with their relative position in the other, but these first $K$ nodes need not be the same on each measure.
@@
From these figures we can see that the rankings from the pairwise projection networks and the ones from our new method are positively correlated (as they should), being very similar in some cases (e.g. subfigures a and b) but different enough that they represent a different meaning of what importance is. The “oscillations” in the intermediate range point to the fact that, even though the correlation tends stabilize when we consider most nodes, there are some intermediately ranked nodes in a measure, considered very relevant in the other, causing the down spikes. We have also briefly noted this in the manuscript.
@@ 4.3 $k$-step hypergraphs
In [xu2023two] the authors designed a new centrality measure (the two-step eigenvector centrality) for undirected graphs, which consisted of constructing a tensor (the two-step tensor) from powers of the adjacency matrix of the graph, which could then leverage the machinery of $\mathcal{H}$-eigenvectors in order to rank the nodes of the original graph [^6] Said article is only concerned with undirected graphs and $\mathcal{H}$-eigenvectors, however there is actually no roadblock for extending it to directed graphs and $\mathcal{Z}$-eigenvectors..
Definition 4.6 (2-step centrality of a graph [xu2023two]).
@@
Definition 4.7 ($k$-step hyperedge).
@@
Clearly, for an undirected network we have the symmetry property $T_{i_{1}i_{2}\dots i_{k-1}i_{k}}=T_{i_{k}i_{k-1}\dots i_{2}i_{1}}$, which is why in [xu2023two] they safely ignore the transposition step, although omitting this discussion.
@@
In some sense this measures provides centrality to the target node from its $k-1$ ancestors. This is very reminiscent of centrality measures involving the iterated line graph or the Hashimoto matrix in standard networks [krzakala2013spectral, aleja2019nonbacktracking].
@@
Theorem 4.8 (Strong connectedness of the hypergraph induced by $T_{i_{1}\dots i_{k}}$).
@@
Proof.
Consider the definition of strong connectivity based on the graph $G^{M}$ (see Definition 2.4)
@@
because if $G$ is strongly connected then there must be at least one $(k-2)$-long walk starting from $j$ (possibly repeating edges and nodes). Therefore $M\geq A$ component-wise, so if $G$ is strongly connected, $G^{M}$ is strongly connected and so is $H$. ∎
@@ 4.3.1 Numerical comparisons
In the last subfigure, corresponding to a flight transportation network, we find it remarkable that the rankings are almost identical, having the Spearman correlation very close to 1 for all $K$ sampled. This last dataset corresponds to an undirected network, which, for the purposes of this article, is treated as a directed one with both directions per edge. However this bears no relation with the high correlation, and indeed we have checked that other undirected networks (e.g. the US Power Grid) do not have this feature.
@@ 5 Conclusions
Our paper sheds light on the wide variety of non-undirected hyperedges, beyond traditional directedness, acknowledging diverse forms of multiple interactions among individuals. By introducing the term “heterogeneous hyperedge”, we aimed to broaden the understanding of non-undirected interactions, including directed hyperedges as a specific type. In order to convey the need of these structures in real data, we listed several real systems which would benefit from such a description; future studies and data collections will surely reveal the relevance in these scenarios. From a more theoretical point of view, several problems are now open for discussion: any prior works relying on directed hypergraphs (e.g. hypergraph curvature [leal2021forman], synchronization [gallo2022synchronization] or community detection [contreras2023detecting]) need to be extended to heterogeneous hypergraphs, specially when adjacency representations are involved.
@@
Funding
This work has been partially supported by projects M3033 (URJC Grant) and the INCIBE/URJC Agreement M3386/2024/0031/001 within the framework of the Recovery, Transformation and Resilience Plan funds of the European Union (Next Generation EU). G. C-A. is funded by the URJC fellowship PREDOC-21-026-2164.
@@
The data used for the numerical results presented in this work, as well as the code written to analyze and plot them, can be found in the repository
[…]